# Causal Structure Learning in Hawkes Processes with Complex Latent Confounder Networks

### v1

@@ Appendix B Multivariate Hawkes Process Details
Before introducing multivariate Hawkes process, we first describe the temporal point process and counting process briefly. A temporal point process is a random process whose realization consists of a list of discrete events in time $\{T_{1},T_{2},\dots\}$ taking values in $0,\infty)$. Another equivalent representation is the counting process, $N_{1}=\{N_{1}(t)|t\in[0,\infty)\}$, where $N_{1}(t)$ records the number of events before time $t$ and $N_{1}(0)=0$. A multivariate point process with $l$ types of events is represented by $l$ counting processes $\{N_{i}\}^{l}_{i=1}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. $N_{i}=\{N_{i}(t)|t\in[0,\infty)\}$, where $N_{i}(t)$ is the number of type-$i$ events occurring before time $t$ and $N_{i}(0)=0$. $\mathbf{U}=\{1,\dots,l\}$ (sometimes abbreviated as $[l]$) represents the set of event types. $\Omega=[0,\infty)\times\mathbf{U}$ is the sample space. $\mathcal{F}=\mathcal{F}(t)$ is a filtration, that is, a non-descreasing family of $\sigma$-algebras which for each time point $t\in\mathbb{R}$, represent the set of event sequences the processes can realize before time $t$. $\mathbb{P}$ is the probability measure. Point processes can be characterized by the conditional intensity function, which models patterns of interest, such as self-triggering or self-correcting behaviors [56]. The conditional intensity function is defined as the expected instantaneous rate of type-$i$ events occurring at time $t$, given the event history:
@@
where $\mathcal{H}(t)=\{(t_{k},i)|t_{k}<t,i\in\mathbf{U}\}$ collects historical events of all types before time t. The multivariate Hawkes process is a class of multivariate point processes characterized by a self-triggering pattern as defined in Definition˜2.1.
@@ Appendix C Identifying Intermediate Latent Subprocesses
Figure 6: Example of an intermediate latent subprocess on the directed path from $O_{2}$ to $O_{1}$. (a) The summary causal graph, where $L_{1}$ is the intermediate latent subprocess. (b) The corresponding window causal graph with two effective lag variables.
As shown in the summary causal graph in Fig. 6(a), $L_{1}$ is an intermediate latent subprocess on the directed path fromthe observed subprocess $O_{2}$ to $O_{3}$. According to Proposition˜3.3, $L_{1}$ is not identifiable and its effect is attributed to $O_{2}$, leading to the inference that $O_{2}$ is the parent cause of $O_{3}$. This is because the influence of $L_{1}$ is indistinguishable from that of $O_{2}$ and can be effectively merged into $O_{2}$.
Consider now the corresponding window causal graph shown in Fig. 6(b). The observed variable set is given by $\mathbf{O}_{v}\coloneq\{O_{i}^{(j)}\}^{j\in\{n-m,\dots,n\}}_{i\in\{1,2,3\}}$, where $m=3$ exceeds the number of effective lag variables (which is 2 in this example). Instead of conditioning on all three lagged variables $\{O_{2}^{(n-1)},O_{2}^{(n-2)},O_{2}^{(n-3)}\}$ of $O_{2}$, we exclude $O_{2}^{(n-1)}$ and condition only on $\{O_{2}^{(n-2)},O_{2}^{(n-3)}\}$. In this case, $O_{3}^{(n)}$ becomes d-separated from the remaining variables in $\mathbf{O}_{v}$. This property arises because, due to the presence of the intermediate latent subprocess $L_{1}$, $O_{2}^{(n-1)}$ no longer has a direct influence on $O_{3}^{(n)}$. The following corollary formalizes a general method for identifying the number of intermediate latent subprocesses that may exist between an observed subprocess and each of its inferred observed parent causes.
@@
In Corollary˜C.1, $O_{1}$ and $O_{2}$ may refer to the same subprocess in cases where Proposition˜3.3 infers that $O_{1}$ has a self-loop. In such cases, Corollary˜C.1 can be used to determine whether this self-loop represents a direct self-excitation or is mediated through intermediate latent subprocesses.
@@
Let $\mathcal{O}_{\mathcal{G}}\coloneqq\{O_{i}\}_{i=1}^{p}$ and $\mathbf{O}_{v}\coloneqq\{O_{i}^{(j)}\}_{i\in\{1,2,\dots,p\}}^{j\in\{n-m,\dots,n\}}$. Consider an observed subprocess $O_{1}$ and its inferred parent cause set $\mathcal{P}_{\mathcal{G}}$. For any $O_{2}\in\mathcal{P}_{\mathcal{G}}$, assume the shortest directed path from $O_{2}$ to $O_{1}$ consists of $h$ latent subprocesses. This implies that the lagged variables $\{O_{2}^{(j)}\}_{j\in\{n-h,\dots,n-1\}}$ do not influence $O_{1}^{(n)}$, while the variables $\{O_{2}^{(j)}\}_{j\in\{n-m,\dots,n-h\}}$ do.
Thus, the variable set $\mathbf{P}_{v}=\{O_{i}^{(j)}\}_{O_{i}\in\mathcal{P}_{\mathcal{G}}}^{j\in\{n-m,\dots,n-1\}}\setminus\{O_{2}^{(j)}\}_{j\in\{n-h,\dots,n-1\}}$ is the minimal set that d-separates $O_{1}^{(n)}$ from the remaining variables. By Lemma˜3.2, this implies:
@@ Appendix D Rank Faithfulness for the Hawkes Process
Condition 2 (Rank Faithfulness for the Hawkes Process).
A probability distribution $p$ is rank faithful to the graph $\mathcal{G}$ if every rank constraint on any sub-covariance matrix that holds in $p$ is entailed by every linear structural model (as defined in Eq. 1) with respect to $\mathcal{G}$ and the excitation function $\phi_{ij}(s)=a_{ij}w(t),\,\forall i,j\in\{1,\dots,l\}$.
The rank faithfulness assumption is widely adopted in the causal discovery literature for i.i.d. data [49, 22]. In our setting, it concerns only the excitation function coefficients $a_{ij}$, and prior studies have shown that violations of this assumption occur only in degenerate cases of Lebesgue measure zero. Specifically, it fails only in rare pathological scenarios, such as when multiple $a_{ij}$ coefficients involving those of latent subprocesses are exactly equal across different subprocesses in a manner that induces rank deficiency—situations that are highly unlikely to arise in practical applications.
To empirically assess the robustness of our method to potential violations of rank faithfulness, we conduct a sensitivity analysis where, for each synthetic graph, we choose the exponential excitation function $\phi_{ij}(s)=\alpha_{ij}e^{-\beta s}$ and deliberately assign identical $a_{ij}$ values to two randomly selected edges, thereby artificially increasing the risk of the violation of rank faithfulness. The results, reported in Table˜2, demonstrate that our method remains robust even under such perturbations.
@@ Appendix E Accounting for Self-Looped Observed Subprocesses under Latent Confounder Influence
Figure 7: Illustration of self-Looped observed subprocesses under latent confounder influence. (a) Summary causal graph where $O_{1},O_{2},O_{3},$ and $O_{4}$ are observed subprocesses, and $L_{1}$ is a latent confounder subprocess. All subprocesses have self-loops. (b) Corresponding window causal graph for (a), illustrating the discretized causal mechanisms among $O_{1},O_{2}$, and $L_{1}$, with two effective lag variables.
@@
Fortunately, since these lagged variables are observed, they can be explicitly incorporated into the structural equations and, correspondingly, into the covariance matrix. Considering the window graph in Fig. 7(b) with $m$ effective lag variables, the structural equations for the observed variables $\{O_{i}^{(j)}\}_{i\in\{1,2\}}^{j\in\{n-m,\dots,n\}}$ can be written as:
@@
where $2m$ corresponds to the observed lagged variables of $O_{1}$ and $O_{2}$, and 1 corresponds to the latent confounder subprocess $L_{1}$. For a formal proof, see Proposition˜3.5 and Appendix˜K. This result implies the presence of a latent confounder subprocess $L_{1}$, such that the set $\{L_{1},O_{1},O_{2}\}$ forms the parent cause set of $\{O_{1},O_{2}\}$. Conditioning on this set renders $\{O_{1},O_{2}\}$ locally independent of $O_{3}$ and $O_{4}$.
Appendix F Proof of Proposition˜2.5
@@
On the other hand, as discussed in the last paragraph of Section˜2.1 and in [14], $N_{i}$ is said to be locally independent of the rest of the system given a minimal set $\mathcal{P}_{\mathcal{G}2}\subseteq\mathcal{N}_{\mathcal{G}}$ if and only if $\int_{0}^{t}\phi_{ij}(t-s)\,dN_{j}(s)>0$ for each $N_{j}\in\mathcal{P}_{\mathcal{G}2}$, and zero otherwise.
Therefore, we can see $\mathcal{P}_{\mathcal{G}1}=\mathcal{P}_{\mathcal{G}2}$. $N_{i}$ is locally independent of $\mathcal{N}_{\mathcal{G}}\backslash\mathcal{P}_{\mathcal{G}}$ given $\mathcal{P}_{\mathcal{G}}$ if and only if $\mathcal{P}_{\mathcal{G}}$ is the parent cause set for $N_{i}$ in $\mathcal{N}_{\mathcal{G}}$. This completes the proof of Proposition˜2.5. ∎
Appendix G Proof of Theorem˜3.1
@@
To prove Theorem˜3.1, we proceed in three steps. First, we define the multivariate INAR sequence (Definition˜G.1) and show that it admits a linear autoregressive model representation (Proposition˜G.3). Then, in Theorem˜G.5, we establish that this multivariate INAR counting process converges weakly to a multivariate Hawkes process as the bin size $\Delta\rightarrow 0$, with the correspondence between the parameters of both models made explicit. The details are as follows:
@@
We begin by introducing the multivariate INAR model, adapted from Definition 20 in the paper B. Hawkes forests in [31].
Definition G.1 (Multivariate integer-valued autoregressive model [31]).
An multivariate integer-valued autoregressive time series(multivariate INAR) is a sequence of $\mathbb{N}_{0}$-valued random variables $\mathbf{X}_{v}=\{X_{1}^{(n)},X_{2}^{(n)},\dots,X_{l}^{(n)}\}_{n\in\mathbb{N}_{0}}$ with $X_{i}^{(0)}=0$, defined as:
$X_{i}^{(n)}=\sum_{j=1}^{l}\sum_{k=1}^{n}\sum_{h=1}^{X^{(n-k)}_{j}}\xi_{h}^{(\theta^{(k)}_{ij})}+\epsilon^{(n)}_{i},\quad i\in\{1,\dots,l\},n\in\mathbb{N}_{0},$ (7)
@@
Definition˜G.1 follows Definition 20 in paper B. Hawkes Forests in [31], but with adapted notation to match Theorem˜3.1. Key correspondences include: $d\rightarrow l$, $i\rightarrow j$, $j\rightarrow i$, $l\rightarrow h$, $(\mathbf{X}_{n})_{n\in\mathbb{Z}}\rightarrow\mathbf{X}_{v}$, $X_{n}^{(j)}\rightarrow X_{i}^{(n)}$, $\xi_{n,l}^{(i,j,k)}\rightarrow\xi_{h}^{(\theta^{(k)}_{ij})}$, $\epsilon_{n}^{(j)}\rightarrow\epsilon_{i}^{(n)}$, $\alpha_{i,j,k}\rightarrow\theta^{(k)}_{ij}$, $\alpha_{0,j}\rightarrow\theta^{(0)}_{i}$. We also restrict indices to $n\in\mathbb{N}_{0}$ to match our Hawkes process formulation (Definition˜2.1); this is purely notational and does not affect the model semantics, as the indices are used to describe relative positions within the time series.
@@
The multivariate INAR sequence admits an equivalent linear autoregressive representation, as shown in Proposition˜G.3, corresponding to Proposition 3.1 in [29]. The current variable $X_{i}^{(n)}$ is expressed as a weighted sum of all lag variables $X_{j}^{n-k}$, plus a constant term $\theta_{i}^{(0)}$ and a stationary white-noise term $\varepsilon_{i}^{(n)}$.
@@
$\varepsilon_{i}^{(n)}\coloneq X_{i}^{(n)}-\theta_{i}^{(0)}-\sum_{j=1}^{l}\sum_{k=1}^{n}\theta_{i,j}^{k}X_{j}^{(n-k)},\quad n\in\mathbb{N}_{0},$ (8)
defines a white-noise sequence, i.e., $(\varepsilon_{i}^{(n)})$ is stationary, $\mathbb{E}[\varepsilon_{i}^{(n)}]=0$, $i\in\{1,\dots,l\}$, $n\in\mathbb{N}_{0}$. Moreover, let the $l\times l$ noise matrices $\mathbf{u}_{n}\mathbf{u}_{n^{\prime}}^{\top}\coloneq[\varepsilon_{i}^{(n)}\varepsilon_{j}^{(n)}]_{(i,j)\in\{1,\dots,l\}}$ and reproduction-coefficient matrices $A_{k}\coloneq[\theta_{ij}^{(k)}]_{(i,j)\in\{1,\dots,l\}}$, we have:
@@
Proposition˜G.3 is adapted from Proposition 3.1 of [29], which also appears as Proposition 6 of the same paper in the author’s doctoral thesis [31]. The original formulation uses full vector and matrix notation; here, we present each dimension separately for consistency with our notation. Moreover, we adapted notations as in Remark˜G.2.
@@
Theorem G.5 (Multivariate INAR converging to multivariate Hawkes process [31]).
Let $\mathcal{N}_{\mathcal{G}1}=\{N_{i}\}^{l}_{i=1}$ be a stationary multivariate Hawkes process with background intensities $\{\mu_{i}\}_{i=1}^{l}$, and piecewise-continuous excitation functions $\{\phi_{ij}(s)\geq 0,\forall s\in(0,\infty)\}_{i=1}^{l}$. For bin width $\Delta\in(0,\delta)$, let $\mathbf{X}_{v}=\{X_{1}^{(n)},X_{2}^{(n)},\dots,X_{l}^{(n)})_{n\in\mathbb{N}_{0}}$ be an multivariate INAR sequence with:
@@
Theorem˜G.5 is a simplified version of Theorem 25 in [31]. The original proof proceeds via convergence of Hawkes forests (constructed via branching random walks), showing that the Hawkes process is a limit of INAR-based approximating forests. The convergence of Hawkes process and INAR comes from the convergence of Hawkes forest and the approximating forest with corresponding parameters. We adapt it here with a direct correspondence between Hawkes and INAR parameters, and restrict domains to $t\in[0,\infty)$ and $n\in\mathbb{N}_{0}$ for consistency and clarification. Typically, Hawkes process results hold for both domains [33, Remark 2], since variable $t$ and $n$ is used only to calibrate relative positions. Moreover, besides the notation changes in Remark˜G.2, we adopt: $\mathbf{N_{F}}\rightarrow\mathcal{N}_{\mathcal{G}1}$, $\mathbf{N_{F^{(\Delta)}}}\rightarrow\mathcal{N}_{\mathcal{G}2}$, the reproduction intensities $h_{i,j}=w_{i,j}m_{i,j}$ $\rightarrow$ excitation function $\phi_{ij}$.
@@
The constant $\delta$ in the Theorem˜G.5 comes from the moment structure of the INAR sequence. For details, see Theorem 2 in [30] and Corollary 24 in paper B. Hawkes forests in [31].
@@
The linear autoregressive representation of the multivariate INAR model is established in Proposition˜G.3, based on the model definition provided in Definition˜G.1. The convergence of the multivariate INAR process to the multivariate Hawkes process, along with the correspondence of their parameters, is presented in Theorem˜G.5. Together, these results validate the discrete-time linear formulation stated in Theorem˜3.1. This completes the proof.
Appendix H Proof of Lemma˜3.2
@@
The proof of Lemma˜3.2 is based on Proposition 2.2 and Theorem 2.4 from [50], which we restate here for completeness.
Proposition H.1 (Rank Characterization of Conditional Independence [50]).
@@
Theorem H.2 (Conditional Independence in Directed Graphical Models [50]).
@@
Since the window causal graph in PO-MHP is a DAG with linear causal relations and serially uncorrelated white noise, the above rank condition applies directly to the window causal graph in the PO-MHP framework. This completes the proof. ∎
Appendix I Proof of Proposition˜3.3

### v2

@@ Appendix B Multivariate Hawkes Process Details
Before introducing multivariate Hawkes process, we first describe the temporal point process and counting process briefly. A temporal point process is a random process whose realization consists of a list of discrete events in time $\{T_{1},T_{2},\dots\}$ taking values in $0,\infty)$. Another equivalent representation is the counting process, $N_{1}=\{N_{1}(t)|t\in[0,\infty)\}$, where $N_{1}(t)$ records the number of events before time $t$ and $N_{1}(0)=0$. A multivariate point process with $l$ types of events is represented by $l$ counting processes $\{N_{i}\}^{l}_{i=1}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. $N_{i}=\{N_{i}(t)|t\in[0,\infty)\}$, where $N_{i}(t)$ is the number of type-$i$ events occurring before time $t$ and $N_{i}(0)=0$. $\mathbf{U}=\{1,\dots,l\}$ (sometimes abbreviated as $[l]$) represents the set of event types. $\Omega=[0,\infty)\times\mathbf{U}$ is the sample space. $\mathcal{F}=\mathcal{F}(t)$ is a filtration, that is, a non-descreasing family of $\sigma$-algebras which for each time point $t\in\mathbb{R}$, represent the set of event sequences the processes can realize before time $t$. $\mathbb{P}$ is the probability measure. Point processes can be characterized by the conditional intensity function, which models patterns of interest, such as self-triggering or self-correcting behaviors [Xu et al., 2015]. The conditional intensity function is defined as the expected instantaneous rate of type-$i$ events occurring at time $t$, given the event history:
@@
where $\mathcal{H}(t)=\{(t_{k},i)|t_{k}<t,i\in\mathbf{U}\}$ collects historical events of all types before time t. The multivariate Hawkes process is a class of multivariate point processes characterized by a self-triggering pattern as defined in Eq. 1.
@@ Appendix C Identifying Intermediate Latent Subprocesses
Figure 6: Example of an intermediate latent subprocess on the directed path from $O_{2}$ to $O_{1}$. (a) The summary causal graph, where $L_{1}$ is the intermediate latent subprocess. (b) The corresponding window causal graph with two effective lags.
As shown in the summary causal graph in Fig. 6(a), $L_{1}$ is an intermediate latent subprocess on the directed path fromthe observed subprocess $O_{2}$ to $O_{3}$. According to Proposition˜4.3, $L_{1}$ is not identifiable and its effect is attributed to $O_{2}$, leading to the inference that $O_{2}$ is the parent cause of $O_{3}$. This is because the influence of $L_{1}$ is indistinguishable from that of $O_{2}$ and can be effectively merged into $O_{2}$.
Consider now the corresponding window causal graph shown in Fig. 6(b). The observed variable set is given by $\mathbf{O}_{v}\coloneq\{O_{i}^{(j)}\}^{j\in\{n-m,\dots,n\}}_{i\in\{1,2,3\}}$, where $m=3$ exceeds the number of effective lags (which is 2 in this example). Instead of conditioning on all three lagged variables $\{O_{2}^{(n-1)},O_{2}^{(n-2)},O_{2}^{(n-3)}\}$ of $O_{2}$, we exclude $O_{2}^{(n-1)}$ and condition only on $\{O_{2}^{(n-2)},O_{2}^{(n-3)}\}$. In this case, $O_{3}^{(n)}$ becomes d-separated from the remaining variables in $\mathbf{O}_{v}$. This property arises because, due to the presence of the intermediate latent subprocess $L_{1}$, $O_{2}^{(n-1)}$ no longer has a direct influence on $O_{3}^{(n)}$. The following corollary formalizes a general method for identifying the number of intermediate latent subprocesses that may exist between an observed subprocess and each of its inferred observed parent causes.
@@
In Corollary˜C.1, $O_{1}$ and $O_{2}$ may refer to the same subprocess in cases where Proposition˜4.3 infers that $O_{1}$ has a self-loop. In such cases, Corollary˜C.1 can be used to determine whether this self-loop represents a direct self-excitation or is mediated through intermediate latent subprocesses.
@@
Let $\mathcal{O}_{\mathcal{G}}\coloneqq\{O_{i}\}_{i=1}^{p}$ and $\mathbf{O}_{v}\coloneqq\{O_{i}^{(j)}\}_{i\in\{1,2,\dots,p\}}^{j\in\{n-m,\dots,n\}}$. Consider an observed subprocess $O_{1}$ and its inferred parent-cause set $\mathcal{P}_{\mathcal{G}}$. For any $O_{2}\in\mathcal{P}_{\mathcal{G}}$, assume the shortest directed path from $O_{2}$ to $O_{1}$ consists of $h$ latent subprocesses. This implies that the lagged variables $\{O_{2}^{(j)}\}_{j\in\{n-h,\dots,n-1\}}$ do not influence $O_{1}^{(n)}$, while the variables $\{O_{2}^{(j)}\}_{j\in\{n-m,\dots,n-h\}}$ do.
Thus, the variable set $\mathbf{P}_{v}=\{O_{i}^{(j)}\}_{O_{i}\in\mathcal{P}_{\mathcal{G}}}^{j\in\{n-m,\dots,n-1\}}\setminus\{O_{2}^{(j)}\}_{j\in\{n-h,\dots,n-1\}}$ is the minimal set that d-separates $O_{1}^{(n)}$ from the remaining variables. By Lemma˜4.2, this implies:
@@ Appendix D Rank Faithfulness for the Hawkes Process
Assumption 2 (Rank Faithfulness for the Hawkes Process).
A probability distribution $p$ is rank faithful to the graph $\mathcal{G}$ if every rank constraint on any sub cross-covariance matrix that holds in $p$ is entailed by every linear structural model (as defined in Eq. 2) with respect to $\mathcal{G}$ and the excitation function $\phi_{ij}(s)=a_{ij}w(t),\,\forall i,j\in\{1,\dots,l\}$.
The rank faithfulness assumption is widely adopted in the causal discovery literature for i.i.d. data [Spirtes, 2013, Huang et al., 2022]. In our setting, it concerns only the excitation function coefficients $a_{ij}$, and prior studies have shown that violations of this assumption occur only in degenerate cases of Lebesgue measure zero. Specifically, it fails only in rare pathological scenarios, such as when multiple $a_{ij}$ coefficients involving those of latent subprocesses are exactly equal across different subprocesses in a manner that induces rank deficiency—situations that are highly unlikely to arise in practical applications.
To empirically assess the robustness of our method to potential violations of rank faithfulness, we conduct a sensitivity analysis where, for each synthetic graph, we choose the exponential excitation function $\phi_{ij}(s)=\alpha_{ij}e^{-\beta s}$ and deliberately assign identical $a_{ij}$ values to two randomly selected edges, thereby artificially increasing the risk of the violation of rank faithfulness. The results, reported in Table˜3 in Section˜Q.3, demonstrate that our method remains robust even under such perturbations.
@@ Appendix E Accounting for Self-Looped Observed Subprocesses under Latent Confounder Influence
Figure 7: Illustration of self-Looped observed subprocesses under latent confounder influence. (a) Summary causal graph where $O_{1},O_{2},O_{3},$ and $O_{4}$ are observed subprocesses, and $L_{1}$ is a latent confounder subprocess. All subprocesses have self-loops. (b) Corresponding window causal graph for (a), illustrating the discretized causal mechanisms among $O_{1},O_{2}$, and $L_{1}$, with two effective lags.
@@
Fortunately, since these lagged variables are observed, they can be explicitly incorporated into the structural equations and, correspondingly, into the covariance matrix. Considering the window graph in Fig. 7(b) with $m$ effective lagged variables, the structural equations for the observed variables $\{O_{i}^{(j)}\}_{i\in\{1,2\}}^{j\in\{n-m,\dots,n\}}$ can be written as:
@@
where $2m$ corresponds to the observed lagged variables of $O_{1}$ and $O_{2}$, and 1 corresponds to the latent confounder subprocess $L_{1}$. For a formal proof, see Proposition˜4.5 and Appendix˜K. This result implies the presence of a latent confounder subprocess $L_{1}$, such that the set $\{L_{1},O_{1},O_{2}\}$ forms the parent-cause set of $\{O_{1},O_{2}\}$. Conditioning on this set renders $\{O_{1},O_{2}\}$ locally independent of $O_{3}$ and $O_{4}$.
Appendix F Proof of Proposition˜3.4
@@
On the other hand, as discussed in the last paragraph of Section˜3.1 and in [Didelez, 2008], $N_{i}$ is said to be locally independent of the rest of the system given a minimal set $\mathcal{P}_{\mathcal{G}2}\subseteq\mathcal{N}_{\mathcal{G}}$ if and only if $\int_{0}^{t}\phi_{ij}(t-s)\,dN_{j}(s)>0$ for each $N_{j}\in\mathcal{P}_{\mathcal{G}2}$, and zero otherwise.
Therefore, we can see $\mathcal{P}_{\mathcal{G}1}=\mathcal{P}_{\mathcal{G}2}$. $N_{i}$ is locally independent of $\mathcal{N}_{\mathcal{G}}\backslash\mathcal{P}_{\mathcal{G}}$ given $\mathcal{P}_{\mathcal{G}}$ if and only if $\mathcal{P}_{\mathcal{G}}$ is the parent-cause set for $N_{i}$ in $\mathcal{N}_{\mathcal{G}}$. This completes the proof of Proposition˜3.4. ∎
Appendix G Proof of Theorem˜4.1
@@
To prove Theorem˜4.1, we proceed in three steps. First, we define the multivariate INAR sequence (Definition˜G.1) and show that it admits a linear autoregressive model representation (Proposition˜G.3). Then, in Theorem˜G.5, we establish that this multivariate INAR counting process converges weakly to a multivariate Hawkes process as the bin size $\Delta\rightarrow 0$, with the correspondence between the parameters of both models made explicit. The details are as follows:
@@
We begin by introducing the multivariate INAR model, adapted from Definition 20 in the paper B. Hawkes forests in [Kirchner, 2017].
Definition G.1 (Multivariate integer-valued autoregressive model [Kirchner, 2017]).
An multivariate integer-valued autoregressive time series(multivariate INAR) is a sequence of $\mathbb{N}_{0}$-valued random variables $\mathbf{X}_{v}=\{X_{1}^{(n)},X_{2}^{(n)},\dots,X_{l}^{(n)}\}_{n\in\mathbb{Z}^{+}}$ with $X_{i}^{(0)}=0$, defined as:
$X_{i}^{(n)}=\sum_{j=1}^{l}\sum_{k=1}^{n}\sum_{h=1}^{X^{(n-k)}_{j}}\xi_{h}^{(\theta^{(k)}_{ij})}+\epsilon^{(n)}_{i},\quad i\in\{1,\dots,l\},n\in\mathbb{Z}^{+},$ (7)
@@
Definition˜G.1 follows Definition 20 in paper B. Hawkes Forests in [Kirchner, 2017], but with adapted notation to match Theorem˜4.1. Key correspondences include: $d\rightarrow l$, $i\rightarrow j$, $j\rightarrow i$, $l\rightarrow h$, $(\mathbf{X}_{n})_{n\in\mathbb{Z}}\rightarrow\mathbf{X}_{v}$, $X_{n}^{(j)}\rightarrow X_{i}^{(n)}$, $\xi_{n,l}^{(i,j,k)}\rightarrow\xi_{h}^{(\theta^{(k)}_{ij})}$, $\epsilon_{n}^{(j)}\rightarrow\epsilon_{i}^{(n)}$, $\alpha_{i,j,k}\rightarrow\theta^{(k)}_{ij}$, $\alpha_{0,j}\rightarrow\theta^{(0)}_{i}$. We also restrict indices to $n\in\mathbb{Z}^{+}$ to match our Hawkes process formulation (Eq. 1); this is purely notational and does not affect the model semantics, as the indices are used to describe relative positions within the time series.
@@
The multivariate INAR sequence admits an equivalent linear autoregressive representation, as shown in Proposition˜G.3, corresponding to Proposition 3.1 in [Kirchner, 2016a]. The current variable $X_{i}^{(n)}$ is expressed as a weighted sum of all lagged variables $X_{j}^{n-k}$, plus a constant term $\theta_{i}^{(0)}$ and a stationary white-noise term $\varepsilon_{i}^{(n)}$.
@@
$\varepsilon_{i}^{(n)}\coloneq X_{i}^{(n)}-\theta_{i}^{(0)}-\sum_{j=1}^{l}\sum_{k=1}^{n}\theta_{i,j}^{k}X_{j}^{(n-k)},\quad n\in\mathbb{Z}^{+},$ (8)
defines a white-noise sequence, i.e., $(\varepsilon_{i}^{(n)})$ is stationary, $\mathbb{E}[\varepsilon_{i}^{(n)}]=0$, $i\in\{1,\dots,l\}$, $n\in\mathbb{Z}^{+}$. Moreover, let the $l\times l$ noise matrices $\mathbf{u}_{n}\mathbf{u}_{n^{\prime}}^{\top}\coloneq[\varepsilon_{i}^{(n)}\varepsilon_{j}^{(n)}]_{(i,j)\in\{1,\dots,l\}}$ and reproduction-coefficient matrices $A_{k}\coloneq[\theta_{ij}^{(k)}]_{(i,j)\in\{1,\dots,l\}}$, we have:
@@
Proposition˜G.3 is adapted from Proposition 3.1 of [Kirchner, 2016a], which also appears as Proposition 6 of the same paper in the author’s doctoral thesis [Kirchner, 2017]. The original formulation uses full vector and matrix notation; here, we present each dimension separately for consistency with our notation. Moreover, we adapted notations as in Remark˜G.2.
@@
Theorem G.5 (Multivariate INAR converging to multivariate Hawkes process [Kirchner, 2017]).
Let $\mathcal{N}_{\mathcal{G}1}=\{N_{i}\}^{l}_{i=1}$ be a stationary multivariate Hawkes process with background intensities $\{\mu_{i}\}_{i=1}^{l}$, and piecewise-continuous excitation functions $\{\phi_{ij}(s)\geq 0,\forall s\in(0,\infty)\}_{i=1}^{l}$. For bin width $\Delta\in(0,\delta)$, let $\mathbf{X}_{v}=\{X_{1}^{(n)},X_{2}^{(n)},\dots,X_{l}^{(n)})_{n\in\mathbb{Z}^{+}}$ be an multivariate INAR sequence with:
@@
Theorem˜G.5 is a simplified version of Theorem 25 in [Kirchner, 2017]. The original proof proceeds via convergence of Hawkes forests (constructed via branching random walks), showing that the Hawkes process is a limit of INAR-based approximating forests. The convergence of Hawkes process and INAR comes from the convergence of Hawkes forest and the approximating forest with corresponding parameters. We adapt it here with a direct correspondence between Hawkes and INAR parameters, and restrict domains to $t\in[0,\infty)$ and $n\in\mathbb{Z}^{+}$ for consistency and clarification. Typically, Hawkes process results hold for both domains [Laub et al., 2015, Remark 2], since variable $t$ and $n$ is used only to calibrate relative positions. Moreover, besides the notation changes in Remark˜G.2, we adopt: $\mathbf{N_{F}}\rightarrow\mathcal{N}_{\mathcal{G}1}$, $\mathbf{N_{F^{(\Delta)}}}\rightarrow\mathcal{N}_{\mathcal{G}2}$, the reproduction intensities $h_{i,j}=w_{i,j}m_{i,j}$ $\rightarrow$ excitation function $\phi_{ij}$.
@@
The constant $\delta$ in the Theorem˜G.5 comes from the moment structure of the INAR sequence. For details, see Theorem 2 in [Kirchner, 2016b] and Corollary 24 in paper B. Hawkes forests in [Kirchner, 2017].
@@
The linear autoregressive representation of the multivariate INAR model is established in Proposition˜G.3, based on the model definition provided in Definition˜G.1. The convergence of the multivariate INAR process to the multivariate Hawkes process, along with the correspondence of their parameters, is presented in Theorem˜G.5. Together, these results validate the discrete-time linear formulation stated in Theorem˜4.1. This completes the proof. ∎
Appendix H Proof of Lemma˜4.2
@@
The proof of Lemma˜4.2 is based on Proposition 2.2 and Theorem 2.4 from [Sullivant et al., 2010], which we restate here for completeness.
Proposition H.1 (Rank Characterization of Conditional Independence [Sullivant et al., 2010]).
@@
Theorem H.2 (Conditional Independence in Directed Graphical Models [Sullivant et al., 2010]).
@@
The window causal graph in PO-MHP is a directed acyclic graph with linear causal relations and serially uncorrelated white noise. Therefore, the above rank condition applies directly to the window causal graph in the PO-MHP framework. This completes the proof. ∎
Appendix I Proof of Proposition˜4.3